OpenStudy (anonymous):

Point A is located at (0, 4) and point B is located at (-2, -3). Find the point that is 1/4 the distance from point A to point B. Select one: a. (-1, 0.75) b. (-0.75, 2) c. (-0.5, 2.25) d. (-0.25, 3)

OpenStudy (kc_kennylau):

Use the distance formula 5 times to find out the distance between AB, Aa, Ab, Ac and Ad? (a,b,c,d are the options)

OpenStudy (kc_kennylau):

wait

OpenStudy (kc_kennylau):

Use the mid-point formula twice

OpenStudy (kc_kennylau):

\[\Large x_m=\frac{x_1+x_2}2\\\Large y_m=\frac{y_1+y_2}2\]

OpenStudy (anonymous):

I think this i dont know wether it is correct or not.... then by appling the ratio.....

OpenStudy (kc_kennylau):